# Disentangled 4D Gaussian Splatting: Towards Faster and More Efficient Dynamic Scene Rendering

### v1

@@ 3.2.2 The Differentiable Projection First Splatting for 4D Gaussian
$\varphi(\mu)=\mathbf{W}_{view}\mu+d_{view}.$ (12)
The velocity of mean in camera space $\mathcal{V}_{view}$ can be obtained as:
$\mathcal{V}_{view}=\varphi^{{}^{\prime}}(\mathcal{V}_{3D}),$ (13)
$\varphi^{{}^{\prime}}(\mathcal{V})=\mathbf{W}_{view}\mathcal{V}$
@@
$\begin{pmatrix}v_{0}\\ v_{1}\\ v_{2}\end{pmatrix}=\phi^{{}^{\prime}}_{3D}(\mathcal{V}_{view})=\begin{pmatrix}\mathcal{V}_{view_{x}}/P_{2}\\ \mathcal{V}_{view_{y}}/P_{2}\\ \left\|(\mathcal{V}_{view_{x}},\mathcal{V}_{view_{y}},\mathcal{V}_{view_{z}})^{T}\right\|\end{pmatrix}.$ (15)
Following 3DGS we apply the Jacobian matrix of $\phi_{3D}$, which can be formulated as:
@@
$\Sigma^{{}^{\prime}}=JW_{view}\Sigma W_{view}^{T}J^{T}.$ (18)
Then we utilize the fast rasterization technique [10] to get the image in the end.
@@
Table 1: Comparison with the state-of-the-art methods on the Plenoptic Video benchmark. *:Only tested on the scene Flame Salmon. **:Results on Spinach,Beef,and Steak scenes.
3.3 Spatial Edge Loss Temporal Split Strategy
Artifacts significantly affect 3DGS [30] rendering quality. Unlike static scenes, dynamic scenes present both spatial and temporal artifact issues. To address the artifact problem in temporal and spatial, we utilize spatial edge loss and temporal split strategy to relieve this. Spatially, we have noticed that artifacts can lead to the abnormal of high-frequency components in image. Therefore we utilize the Laplacian convolution to obtain high frequency part and calculate L1 loss specifically for high-frequency components. It can be defined as:
$L_{edge}=Loss_{l1}(image_{L}^{gt},image_{L}^{render})$ (19)
where $image_{L}^{gt},image_{L}^{render}$ denote ground truth and rendered images post-Laplacian convolution. Our total loss can be formulated as:
$L=(1-\lambda_{1})L_{1}+\lambda_{1}L_{ssim}+\lambda_{2}L_{edge}$ (20)
Temporally, to handle the inconsistencies in the spatiotemporal frequency domain and avoid the overconfidence in time dimension we decouple the traditional splitting strategy based on $xyz$ coordinate gradient threshold into two strategies: time-domain splitting and spatial splitting. And the time-domain splitting based on the t coordinate gradient, spatial splitting based on the $xyz$ coordinate gradient. The spatial and temporal scale of 4D Gaussians also split separately.
4 Experiments
4.1 Datasets
@@
T-NeRF[20] 29.51 0.95 0.08 - -
D-NeRF[20] 29.67 0.95 0.07 - 33.23
TiNeuVox[7] 32.67 0.97 0.04 1.60 19.46
HexPlanes[4] 31.04 0.97 - - -
K-Planes-explicit[8] 31.05 0.97 - - 25.37
K-Planes-hybrid[8] 31.61 0.97 0.02 - -
V4D[9] 33.72 0.98 0.03 2.08 -
Deformable3DGS[27] 39.31 0.99 0.01 85.45 34.09
RealTime4DGS[28] 32.71 0.97 0.03 289.07 20.80
Rotor4DGS[6] 34.26 0.97 0.03 1257.63 25.48
Ours 33.34 0.97 0.03 1549.03 14.76
Table 2: Evaluation on D-NeRF Dataset
@@
In this section, we evaluate our method on two commonly used datasets, Plenoptic Video Dataset [14] and D-NeRF dataset [20], representative of various challenges in dynamic scene modeling.
@@
Plenoptic Video Dataset [14] contains real-world multiview videos of 6 scenes, each lasting ten seconds. These scenes incorporate complex motions and materials that are reflective or transparent. Following prior work [28], we utilize the colored point cloud generated by COLMAP from the first frame of each scene. And the resolution of $1352\times 1014$ is adopted.
@@
D-NeRF [20] is a monocular video dataset comprising eight videos of synthetic scenes. Following prior work [28], we utilize randomly selected points, evenly distributed within the cubic volume defined by $[-1.2,1.2]^{3}$, and set their initial mean as the time duration of scene.
4.2 Implementation Details
@@
The densification gradient threshold is set as $55e-$ 5 in D-NeRF and $52e-$ 5 in Plenoptic datasets. The rotors are initialized with (1,0,0,0) equivalent of static identity transformation. Learning rates, densification, pruning, and opacity reset settings all follow [10]. Optimizer is Adam following prior work. For each scene we train for 20,000 steps. The LPIPS [31] in the Plenoptic Video dataset and the D-NeRF dataset are computed using AlexNet [11] and VGGNet [21], respectively.
@@
4.3 Results of Dynamic Novel View Synthesis
4.3.1 Results on the multi-view real scenes
As detailed in Tab.1, our approach not only significantly surpasses previous methods in terms of rendering quality but also achieves substantial speed improvements. Our method renders high-resolution videos ($1352\times 1014$) at 343 FPS on an NVIDIA RTX 3090 GPU, which markedly outperforms NeRF based method. And our method also outperforms the slicing first 4D Gaussian [26, 6, 28] on rendering speed and quality. We believe our improvements over [28] mainly come from the new presentation and rendering process we proposed. As Fig.2 illustrated,
4.3.2 Results on the monocular synthetic videos
@@
Monocular video novel view synthesis for dynamic scene is particularly challenging due to its sparse input views[6]. As summarized in Tab.2,Wu et al. [26] achieves the highest PSNR due to its directly racking 3D Gaussian deformation, which perfectly aligns with D-NeRF dataset but can not handle sudden appearances or disappearances very well. Our method achieves the fastest rendering speed among all the methods, and obtain a balanced quality. And our method can handle all kinds of scenes well, unlike [6, 27] may not perform stably in certain scenarios, such as Lego scene. The PSNR of Deformable3DGS [27] and Rotor4DGS [6] are 25.38, 24.93 respectively which are far away from their mean. And in contrast our method is 26.60. The variance of PSNR also demonstrates the stability of our method. The PSNR performance of each scene is shown in supplementary material.
4.3.3 Speed Result of Same Rendering Condition
@@
Table 3: Comparison of 100,000 Gaussians CUDA rendering time
To demonstrate the excellent speed advantage of our method, we test the CUDA rendering time of the traditional 4D Gaussian methods and our method under the same viewing angle and the number of Gaussians is 100,000. The result in Tab. 3 shows that our rendering speed has improved by approximately 26% compared to traditional methods.
4.4 Ablation and Analysis
ID Ablation Items D-NeRF
@@ 4.4 Ablation and Analysis
a 32.89 0.95
b 33.07 0.96
c 33.20 0.96
Full 33.34 0.97
Table 4: Ablation on D-NeRF Dataset. We validate two designs on rendering quality
Figure 3: Comparison of rendering with and without spatial edge loss and time splitting strategy
Tab.4 and Fig.3 ablates the effectiveness of individual designs in our method.
4.4.1 Spatial Edge Loss
@@
As Tab.4 (c) shown, adding spatial edge loss can improve the rendering quality both in PSNR and SSIM. The effect of spatial edge loss is clearly revealed in Fig.3. For example, the detail edge around the scene Lego, Bouncing Balls and T-Rex is more refined. The burrs caused by artifacts are significant reduced. This demonstrates that the spatial edge loss helps reduce the artifacts and improve the expression details. Therefore, our method can handle the complex scenes better like Lego in D-NeRF and Plenoptic Video Dataset.
4.4.2 Temporal Splitting strategy

### v3

@@ III-B2 The Differentiable ”Projection-First” Splatting for 4D Gaussian
$\varphi(\mu)=\mathbf{W}_{\text{view}}\mu+d_{\text{view}}.$ (10)
Second, the velocity of mean in camera space $\mathcal{V}_{view}$ can be obtained as:
$\mathcal{V}_{\text{view }}=\varphi^{{}^{\prime}}(\mathcal{V}_{3D}),$ (11)
$\varphi^{{}^{\prime}}(\mathcal{V})=\mathbf{W}_{\text{view}}\mathcal{V}$
@@
$\begin{pmatrix}v_{0}\\ v_{1}\\ v_{2}\end{pmatrix}=\phi^{{}^{\prime}}_{3D}(\mathcal{V}_{\text{view}})=\begin{pmatrix}\mathcal{V}_{\text{view}_{x}}/P_{2}\\ \mathcal{V}_{\text{view}_{y}}/P_{2}\\ \left\|(\mathcal{V}_{\text{view}_{x}},\mathcal{V}_{\text{view}_{y}},\mathcal{V}_{\text{view}_{z}})^{T}\right\|\end{pmatrix}$. (13)
At last, to derive the covariance matrix $\Sigma^{{}^{\prime}}$ in ray space, we compute the Jacobian matrix $\mathbf{J}_{k}$ of $\phi_{3D}$. Following 3DGS, Jk can be formulated as:
@@
$\Sigma^{{}^{\prime}}=JW_{\text{view}}\Sigma W_{\text{view}}^{T}J^{T}.$ (16)
Then we utilize the fast rasterization technique [17] to get the image in the end. An algorithm outlining the process is provided as algorithm 1.
1Inputs: scales $\mathbf{s}_{4D}=\{\mathbf{s}_{x},\mathbf{s}_{y},\mathbf{s}_{z},\mathbf{s}_{t}\}$,
2 mean $\mu_{4D}=\{\mu_{x},\mu_{y},\mu_{z},\mu_{t}\}$, quaternion $q=\{q_{1},q_{2},q_{3},q_{4}\}$,
3 velocity of mean $\mathcal{V}_{3D}=\{v_{x},v_{y},v_{z}\}$,
4 camera position $\mathbf{P}_{camera}$, camera rotation $\mathbf{R}_{camera}$,
5 timestamp $t_{0}$
6 Outputs: rendered image $I_{output}\in R^{H\times W\times 3}$
7 Initialization:
8 $\mathbf{P}_{camera0}\leftarrow None$, $\mathbf{R}_{camera0}\leftarrow None$
9 while Inputs != None do
10 if $\mathbf{P}_{camera}$ != $\mathbf{P}_{camera0}$ and $\mathbf{R}_{camera}$ != $\mathbf{R}_{camera0}$ then
11    $\mathbf{S}_{3D}\leftarrow diag(\mathbf{s}_{x},\mathbf{s}_{y},\mathbf{s}_{z})$,
12    $\mathbf{R}_{3D}\leftarrow q$,
13    $\Sigma_{3D}\leftarrow\mathbf{R}_{3D}\mathbf{S}_{3D}\mathbf{S}_{3D}^{T}\mathbf{R}_{3D}^{T}$,
14    $\triangleright$ calculate 3D covariance matrix $\Sigma_{3D}$
15    $\mathbf{P}_{view}\leftarrow project(\mathbf{P}_{camera}$$\mathbf{P}_{camera},\mu_{3d})$
16    $\triangleright$ calculate mean in ray space
17    $\mathcal{V}_{view}\leftarrow project(\mathbf{P}_{camera},\mathbf{P}_{camera},\mu_{3d})$
18    $\triangleright$ calculate velocity mean in ray space
19
20   end if
21 $\mathbf{P}_{t_{0}}\leftarrow\mu_{3d}+\mathcal{V}_{3D}\times(\mu_{t}-t_{0})$
22 $\triangleright$ calculate mean in camera space at $t_{0}$
23 $\mathbf{P}_{view,t_{0}}\leftarrow\mathbf{P}_{view}+\mathcal{V}_{view}\times(\mu_{t}-t_{0})$
24 $\triangleright$ calculate mean in ray space at $t_{0}$
25 $\Sigma^{{}^{\prime}}\leftarrow project(\Sigma_{3D},P_{t_{0}})$
26 $\triangleright$ approximate projection of covariance matrix
27 $I_{output}\leftarrow fast\_rasterization(\mathbf{P}_{view,t_{0}},\Sigma^{{}^{\prime}})$
28   output $I_{output}$
29
30 end while
Algorithm 1 Differentiable ”projection-first” splatting for 4D Gaussian.
TABLE I: Comparison with the state-of-the-art methods on the Plenoptic Video benchmark. *: Only tested on the scene Flame Salmon. **: Results on Spinach,Beef,and Steak scenes. Red denotes SOTA. Blue denotes second. Yellow denotes third.
@@ 27 $I_{output}\leftarrow fast\_rasterization(\mathbf{P}_{view,t_{0}},\Sigma^{{}^{\prime}})$
TABLE II: Evaluation on Google Immersive Dataset. “Size/Fr” stands for model size per frame. Red denotes SOTA. Blue denotes second. Yellow denotes third.
[…]
NeRFPlayer[29] 25.8 0.196 0.076 17.1 MB 0.12
HyperReel[29] 28.8 0.193 0.063 1.2 MB 4
SpacetimeGS[7] 29.2 0.081 0.042 1.2 MB 99
Ours 30.6 0.104 0.041 0.95 MB 183
III-C Flow-Gradient Guided Consistency Loss and Temporal Splitting Strategy
Beyond optimizing rendering efficiency, we also aim to enhance rendering quality and ensure coherent object motions in dynamic scenes. Artifacts and distortions, which already degrade 3DGS rendering quality [18], are even more problematic in dynamic settings where inaccurate motion boundaries or inconsistent trajectories may occur. To tackle these challenges, we introduce a flow-gradient guided consistency loss and a temporal splitting strategy.
III-C1 Flow-Gradient Guided Consistency Loss
Most existing dynamic scene methods only supervise the rendered appearance, without explicitly constraining the motion of objects. This often leads to distorted or unstable trajectories, especially around motion boundaries. Since both the color image and optical flow are rendered outputs in our framework, they can be easily obtained without requiring additional annotations. We introduce a regularization to enforce structural consistency between them.
Concretely, we regard image gradients as reliable indicators of structural discontinuities. We first compute the gradient magnitude of the rendered optical flow field:
$M=\sqrt{u^{2}+v^{2}+\varepsilon},$ (17)
and extract the gradient of the RGB image $I$ using a Sobel operator. We then design a consistency loss to encourage flow discontinuities to align with image edges:
$L_{\text{fg}}=\lambda_{flow}\times(\frac{1}{N}\sum_{x,y}\|\nabla M(x,y)\|\cdot\big(1-\|\nabla I(x,y)\|\big)),$ (18)
where $\nabla M(x,y)$ denotes the normalized gradient of the flow magnitude, $\nabla I(x,y)$ is the normalized image gradient and $\lambda_{flow}$ is the scale of flow-gradient guided consistency loss.
This loss penalizes strong flow gradients in regions without corresponding image edges, while allowing sharp motion changes at true structural boundaries. By doing so, the flow-gradient guided consistency loss improves the coherence of rendered motion and produces more faithful dynamic trajectories, thereby enhancing the realism of novel view synthesis in dynamic scenes.
III-C2 Temporal Splitting Strategy
In addition to spatial constraints, dynamic scenes often involve complex, non-linear object motions such as sudden appearance, disappearance, or occlusion changes. Conventional 4DGS approaches approximate object motions with first- or second-order models, which is insufficient for capturing such dynamics. To address this, we decouple the splitting strategy into temporal and spatial components. Specifically, time-domain splitting is guided by the gradient of the $t$ coordinate, while spatial splitting is based on the gradients of xyzxyz coordinates. Moreover, the spatial and temporal scales of 4D Gaussians are split independently, allowing finer-grained modeling of both spatial structures and temporal variations.
IV Experiments
IV-A Datasets
(a) Comparison with the state-of-the-art on flame steak scene
(b) Comparison with the state-of-the-art approaches on coffee martini scene
Figure 3: Visual comparisons on Plenoptic Video Dataset
(a) SpacetimeGS[7]
(b) Ours
(c) Ground Truth
Figure 4: Visual comparisons on Google Immersive Dataset
TABLE III: Evaluation on HyperNeRF Dataset. Red denotes SOTA. Blue denotes second. Yellow denotes third.
Method Chicken Banana Broom 3D Printer Avg
Nerfies[29] 26.7 22.4 19.2 20.6 22.2
HyperNeRF[29] 26.9 23.3 19.3 20.0 22.4
TiNeuVox-B[44] 28.3 24.4 21.5 22.8 24.3
FFDNeRF[38] 28.0 24.3 21.9 22.8 24.2
3D-GS[39] 19.7 20.4 20.6 18.3 19.7
Deformable4DGS[39] 28.7 28.0 22.0 22.1 25.2
Ours 29.4 29.1 22.4 22.3 25.8
TABLE IV: Evaluation on D-NeRF Dataset. Red denotes SOTA. Blue denotes second. Yellow denotes third.
@@ IV-A Datasets
T-NeRF[29] 29.51 0.95 0.08 -
D-NeRF[29] 29.67 0.95 0.07 -
TiNeuVox[44] 32.67 0.97 0.04 1.60
HexPlanes[38] 31.04 0.97 - -
K-Planes-explicit[39] 31.05 0.97 - -
K-Planes-hybrid[39] 31.61 0.97 0.02 -
V4D[45] 33.72 0.98 0.03 2.08
RealTime4DGS[1] 32.71 0.97 0.03 289.07
Rotor4DGS[5] 34.26 0.97 0.03 1257.63
Ours 33.61 0.98 0.02 1549.03
(a) RealTime4DGS[1]
(b) Rotor4DGS[5]
(c) Ours
(d) Ground Truth
Figure 5: Visualization of Chicken scene on HyperNeRF Dataset
(a) RealTime4DGS[1]
(b) Rotor4DGS[5]
(c) Ours
(d) Ground Truth
Figure 6: Visualization of Banana scene on HyperNeRF Dataset
(a) RealTime4DGS[1]
(b) Rotor4DGS[5]
(c) Ours
(d) Ground Truth
Figure 7: Visualization of Lego scene
@@
In this section, we evaluate our method on four commonly used datasets, Plenoptic Video Dataset [2], Google Immersive Dataset[46], HyperNeRF Dataset[47] and D-NeRF dataset [29].
@@
Plenoptic Video Dataset[2] contains real-world multiview videos of 6 scenes, each lasting ten seconds. These scenes include complex motions and materials that are reflective or transparent. Following prior work [1], we utilize the colored point cloud generated by COLMAP from the first frame of each scene. And the resolution of $1352\times 1014$ is adopted.
Google Immersive Dataset[46] contains both indoor and outdoor scenes captured with a $46$-camera rig. The cameras are equipped with fisheye lenses and mounted on an outward-facing hemisphere, which results in less view overlap compared to traditional outside-in setups, thereby posing additional challenges. Following prior work [43, 42], we evaluate on $7$ selected scenes (Welder, Flames, Truck, Exhibit, FacePaint1, FacePaint2, Cave) and hold out the center camera as the test view.
HyperNeRF Dataset[47] contains dynamic real-world scenes exhibiting complex non-rigid deformations and topological changes. The data is captured using one or two handheld cameras with relatively straightforward camera motion. Following prior work [47], we utilize $200$ randomly selected frames from each sequence for training and evaluation.
@@
D-NeRF[29] is a monocular video dataset comprising eight videos of synthetic scenes. Following prior work [1], we utilize randomly selected points, evenly distributed within the cubic volume defined by $1.2,1.2]^{3}$, and set their initial mean as the time duration of scene.
IV-B Implementation Details
@@
The densification gradient threshold is set as 5e55e-5 in D-NeRF and 2e52e-5 in Plenoptic datasets. The rotors are initialized with (1,0,0,0) equivalent of static identity transformation. Learning rates, densification, pruning, and opacity reset settings all follow [17]. Optimizer is Adam following prior work. For each scene we train for 20,000 steps. The LPIPS [48] in the Plenoptic Video dataset and the Google Immersive Dataset[46] are computed using AlexNet [49]. The D-NeRF [29] dataset is computed using VGGNet [50], respectively. To ensure a fair comparison with previous works we do not fix the viewpoints.
@@
IV-C Comparison with Existing Works on Dynamic Novel View Synthesis
IV-C1 Results on the multi-view real-world scenes
Plenoptic Video Dataset [2]
As summarized in Tab. I, our approach substantially surpasses previous methods in both rendering quality and computational efficiency on the Plenoptic Video Dataset [2]. Specifically, our method renders high-resolution videos ($1352\times 1014$) at 343 FPS on an NVIDIA RTX 3090 GPU, representing a significant improvement over all existing approaches. Notably, even when compared with recent slicing-first 4D Gaussian methods [4, 5, 1], our technique exhibits superior rendering fidelity and speed. The comparison with [1] further highlights the efficiency and visual quality advantages of our Disentangled4DGS framework. Visual results in Fig. 3 demonstrate that our method preserves finer dynamic and static details—such as the spray gun in flame steak and the glass cups in coffee martini—surpassing prior methods in both realism and consistency.
Google Immersive Dataset [46]
As presented in Tab. II, we evaluate PSNR, LPIPS, DSSIM, model size per frame, and FPS. Our method achieves the best overall performance across all metrics on a single NVIDIA A6000 GPU under the same experimental settings as [7]: the highest PSNR (30.6), the lowest DSSIM (0.041), the most compact model size (0.95 MB/frame), and the fastest rendering speed (194 FPS). Compared with SpacetimeGS [7], our approach improves PSNR by 1.4 dB while nearly doubling the rendering speed, demonstrating an excellent trade-off among quality, compactness, and efficiency. As shown in Fig. 4, our method also produces fewer visual artifacts, further confirming its robustness in complex immersive scenes.
HyperNeRF Dataset [47]
Following prior works, we report PSNR values on four challenging dynamic scenes (Chicken, Banana, Broom, 3D Printer) in the HyperNeRF dataset (Tab. III). Our method consistently outperforms all baselines, achieving the best results on three out of four scenes and an average PSNR of 25.8, surpassing Deformable4DGS [39] by +0.6 dB on average. Despite comparable performance on 3D Printer, our approach exhibits stronger generalization and motion modeling capabilities, particularly for complex dynamic regions—such as the moving chicken toy in Chicken and the hand motion in Banana—as illustrated in Fig. 5 and Fig. 6.
IV-C2 Results on the monocular synthetic videos
@@
Monocular video novel view synthesis for dynamic scenes remains particularly challenging due to the sparsity of input views [5]. As summarized in Tab.IV, our method achieves the fastest rendering speed among all evaluated techniques, while also delivering a balanced level of quality. A key advantage of our approach is its consistent performance across all types of scenes. This contrasts with methods like Rotor4DGS [5], which can exhibit unstable performance in certain scenarios, such as the Lego scene as shown in Fig. 7. In such cases, the PSNR (24.93) is significantly lower than our method’s 26.60, demonstrating our superior stability. The variance in PSNR across scenes further confirms the robustness of our method. Detailed PSNR performance for each scene is presented in the supplementary material.
IV-C3 Speed Result of Same Rendering Condition
TABLE V: Rendering time comparison on 100,000 Gaussians
@@ IV-C3 Speed Result of Same Rendering Condition
TABLE VI: Ablation studies. We validate two designs on rendering quality on D-NeRF Dataset.
ID Ablation Items D-NeRF Plenoptic Dataset
Edge
Loss
Temporal
Split
@@
(a) 33.20 0.95 32.44 0.013
(b) 33.47 0.97 32.58 0.012
(c) 33.40 0.97 32.56 0.013
Full 33.61 0.98 32.75 0.011
(a) w/ Temporal Splitting
& Edge Loss
(b) w/o Spatial Edge
Loss
(c) w/o Temporal Splitting
strategy
(d) w/o Temporal Splitting
& Edge Loss
Figure 8: Visualization of Lego scene
To demonstrate the rendering speed advantage of our method, we test the CUDA rendering time of the traditional 4D Gaussian methods and our method under under identical conditions, i.e., the same viewing angle, 100,000 Gaussians. The result in Tab. V shows that our approach significantly improves the rendering speed by 25% compared to the most recent method [5]. And When our method deals with scenes with fixed viewpoints, it can simultaneously render multiple frames in parallel, making our method more efficient in terms of speed. Our method consumes an average of 500 MB VRAM on this dataset for each additional frame during parallel rendering, while the delay only increases about 0.4 ms
IV-D Ablation and Analysis
In Tab. III and Fig. 8, we conduct ablation studies on effectiveness of individual designs in our method.
IV-D1 Flow-Gradient Guided Consistency Loss
@@
As shown in Tab. III (c), our flow-gradient guided consistency loss improves the rendering quality in both PSNR and SSIM and the visual effect of the flow-gradient guided consistency loss is clearly illustrated in Fig. 8. Specifically, the details in the edge around the scene Lego, Standup, and Hellwarrior are consistently enhanced while the blurs are significantly reduced with our flow-gradient guided consistency loss. This demonstrates that the flow-gradient guided consistency loss helps reduce the artifacts and improve the high-frequency details. Therefore, our method improves the rendering quality especially for the complex scenes like Lego in D-NeRF and Plenoptic Video Dataset.
IV-D2 Temporal Splitting strategy